Which of the following would eliminate the variable on the left side of the given equation?


36 - 19w = -6w + 41

Answer by Theo(13342)

start with 3 - 19w = -6w + 41
add 19w to both sides of the equation to get:
36 - 19w + 19w = -6w + 41 + 19w
combine like terms to ge:
36 = 13w + 41
you have just eliminated the variable from the left side of the equation.
subtract 41 from both sides of the equation to get:
36 - 41 = 13w + 41 - 41
combine like terms to get:
-5 = 13w
you have just eliminated all constants from the right side of the equqtion.
divide both sides of the equation by 13 to get:
-5/13 = 13w/13
simplify to get:
-5/13 = w
you have just solved for w.
replace w with -5/13 and evaluate the original equation.
you will get:
36 - 19 * -5/13 = -6 * -5/13 + 41
multiply both sides of this equation by 13 to get:
36*13 - 19 * -5 = -6 * -5 + 41 * 13
simplify to get:
468 + 95 = 30 + 533
combine like terms to get:
563 = 563
this confirms the solution is correct.
this took it a lot further than you needed to go.
your solution to the problem as stated would be:
add 192 to both sides of the equation.
that removes the variable from the left side of the equation and moves it to the right side of the equation.